a 7% semiannual treasury bond matures in 10 years and has a par value of $1000. its ytm is 11%. if ytm remains the same, what will the capital gain yield on this bond be over one year?
Answer #1

Semi annual coupon =7%*1000/2 =35
Number of Periods =2*10 =20
Semi annual YTM =11%/2 =5.5%
Price of Bond =PV of Coupons+PV of Par Value =35*((1-(1+5.5%)^-20)/5.5%)+1000/(1+5.5%)^20 =760.9923

Price of Bond After 1 year =PV of Coupons+PV of Par Value =35*((1-(1+5.5%)^-18)/5.5%)+1000/(1+5.5%)^18 =775.0785

Capital gain Yield =(775.0785-760.9923)/760.9923 =1.85%